If you want to double the area of a square, by what percent should you increase the length of its sides.
Answer by CharlesG2(834)

If you want to double the area of a square, by what percent should you increase the length of its sides.
A area square = (s side)^2 --> A = s^2
2 * A = 2 * s^2
sqrt(2s^2) = sqrt(2) * s
sqrt(2) is approximately 1.414
you should increase the length of the sides approx 41.4 %